OpenStudy (anonymous):

Which of the following represents inverse variation? y equals three minus x y equals two divided by x y divided by x equals negative five y equals negative two-thirds x

OpenStudy (anonymous):

im sorry can you please explain in more depth these problems really confuse me

OpenStudy (anonymous):

wait would it be the second option

OpenStudy (radar):

Choose the one where if x should increase, y would decrease. Thats inverse relationahip.

OpenStudy (radar):

That is one if x increases y would decrease.

OpenStudy (triciaal):

inverse variation as one variable increases the other variable decreases and when you multiply the product is always the same. let the variables be x and y then x*y = a number that never changes, k

OpenStudy (anonymous):

if that is so then it is the second option right

OpenStudy (triciaal):

an example let x = 8 and y = 3 then k = 24 if you change and let x = 2 then y = 24/2 = 12 as x decrease then y increase 8 * 3 = 24 2* 12 = 24

OpenStudy (triciaal):

the value of x* y is still 24

OpenStudy (triciaal):

yes it would be the second option

OpenStudy (triciaal):

x*y = 2

OpenStudy (anonymous):

YAY THANK YOU SO MUCH FOR UR HELP opps sorry for the caps

OpenStudy (radar):